# Fedor V. Fomin

### v1

@@ 1 Introduction
Related Work. The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms. Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved. Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16]. This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, ${}^{+}$20].
This progress also concerns structural parameterized complexity, especially in graphs of bounded treewidth [${}^{+}$11, BCKN13, CKN13, FLPS16, BST23]. However, the success of treewidth parameterization does not extend to more general graph width parameters. For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10]. For a comprehensive overview and further details on the algorithms for finding long cycles and paths in graphs, the reader may refer to Chapters 10, 11, and 12 of the book by Cygan et al. [${}^{+}$15].
The interplay of the independence number, connectivity, Hamiltonicity, and path cover is a popular theme in Graph Theory, with a historical foundation dating back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72]. An excellent introduction to this aspect of Graph Theory can be found in the survey by Jackson and Ordaz [JO90]. It is surprising that in graph algorithms, despite tremendous success in the study of Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth, which are sparse graphs, there has been almost no progress on graphs of bounded independence number. The notable exceptions are the works of Ovetsky Fradkin and Seymour [FS15] and Jedlicková and Kratochvíl [JK23]. The first work deals with the edge-disjoint path problem, and the second, most relevant to our first theorem, focuses on Graph Hamiltonicity and $-\ell$-Linkage problems. In the end of this section, we explain the most crucial differences between the approaches taken in our work and the work of Jedlicková and Kratochvíl. Both papers by Ovetsky Fradkin and Seymour and Jedlicková and Kratochvíl provide XP algorithms that run in polynomial time when the independence number of a graph is constant.
Theorem 1 could also be treated as a contribution to the rapidly developing area of algorithms on graphs, excluding some induced subgraphs like $P_{t}$ or $C_{t}$ [${}^{+}$21, ${}^{+}$21]. (Every graph with an independence number of $k$ excludes the graph $kK_{1}$ as an induced subgraph.) However, all results in this area concern the computation of the maximum independent set or, more generally, induced subgraphs of bounded treewidth. Hamiltonicity is a more complex problem; it becomes NP-complete already on $P_{5}$-free graphs (a superclass of split graphs) [BLS99].
Theorem 2 contributes to another popular trend in parameterized algorithms that explores the algorithmic extensions of classic theorems from extremal graph theory. Examples of such works include the algorithmic extensions of Dirac’s theorem by Fomin, Golovach, Sagunov, and Simonov [FGSS20] or Hajnal–Szemerédi’s theorem by Gan, Han and Hu [GHH23].
1.1 Proof overview
Theorem 1: We explain here the main steps of the proof only for $-\ell$-linkages. An extension to arbitrary list TM-embeddings is more technical but builds on similar ideas.
Let us start by sketching the $\operatorname{{\sf X}P}$ algorithm of Jedlicková and Kratochvíl [JK23]. Their algorithm solves $\ell$-Linkage and runs in $n^{f(\alpha(G)+\ell)}$ time. Since Theorem 1 is a strict improvement of the $\operatorname{{\sf X}P}$ algorithm, we highlight crucial differences that allow us to achieve fixed-parameter tractability for the more general problem. An essential step in the algorithm of Jedlicková and Kratochvíl is the following lemma: If a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked, meaning that a path cover of size $\ell$ exists with any given endpoints of the paths.
The algorithm of [JK23] is recursive and requires $k\geq\alpha(G)$ to be given in the input. (This requirement is not a problem for an $\operatorname{{\sf X}P}$ algorithm because $k$ could be computed by a brute force in time $\mathcal{O}(n^{\alpha(G)})$.) When the input graph $G$ is $\Omega(k\cdot\ell)$-connected, the algorithm correctly reports that $G$ is $-\ell$-linked. Otherwise, when the connectivity is smaller, a minimum vertex separator $X$ of $G$ has $\mathcal{O}(k\cdot\ell)$ vertices. Every connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component. For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$. For every connected component of $G-X$ and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component. Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices. By treating each set of entry and exit points as new terminal vertices, the algorithm is called recursively on each component. By restricting the depth of the recursion by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths.
Since the approach of Jedlicková and Kratochvíl exploits small separators separating highly connected components, the first natural direction to explore is whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm. The recursive understanding technique introduced by Chitnis et al. [${}^{+}$16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al. [LRSZ18]. Unfortunately, recursive understanding is not of much help here due to the following reasons. To apply recursive understanding, one has to be able to deal with the primary case of so-called $(p,q)$-unbreakable graphs. However, for our problem, this primary case is not more straightforward than the case of general graphs.
This is why we employ a different method of decomposing the graph. To prevent the $\operatorname{{\sf X}P}$ running time, our algorithm avoids recursive calls to itself. Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either output an independent set of $G$ of size $k$ or find a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of $G-S$ is $\Omega(k\cdot(\ell+|S|))$-connected. Set $S$ breaks solution paths into at most $\ell+2|S|$ parts. Some parts belong to $G[S]$, while others belong to $G-S$. Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time. However, guessing the entry and exit points of these parts, how it is done by Jedlicková and Kratochvíl, would be too time-consuming. Here is precisely the moment when the property of the decomposition comes into play. Each connected component of $G-S$ is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked. Therefore, for any selection of the paths’ endpoints, the required collection of vertex-disjoint paths spanning all component vertices always exists (unless empty). Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of $G-S$ without enumerating the entry-exit endpoints in $G-S$. The latter task reduces to computing a maximum matching in an auxiliary graph. This way, the whole enumeration can be done in $\operatorname{{\sf FPT}}$ time.
Another complication comes from the fact that we do not require the $\alpha(G)<k$ assumption and instead aim to find a solution or an independent set of size $k$. For that, a constructive version of the high connectivity case is required. That is, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we need to be able to construct a family of $\ell$ disjoint paths with the given endpoints spanning the graph or to find an independent set of size $k$. We employ the general result of Thomas and Wollan [TW05]: In a $10\ell$-connected graph, a collection of $\ell$ vertex-disjoint paths with the given endpoints always exists. However, the proof of Thomas and Wollan [TW05] is involved, and it is not apparent whether it could be made constructive. Instead, we employ the property that our graphs are of bounded independence number. We use a Ramsey-based argument, finding either an independent set of size $k$ or a large clique, and in the latter case, we can find the desired $\ell$ paths by connecting the endpoints to the clique via Menger’s theorem.
We remark here that the relation between connectivity and linkages in graphs is well-studied. In particular, a sequence of results, including the above-mentioned work of Thomas and Wollan [TW05], guarantees the existence of $\ell$ vertex-disjoint paths between given pairs of endpoints in a graph of connectivity $\Omega(\ell)$. However, there is a lack of simple and efficient algorithms that compute vertex-disjoint paths connecting the given $\ell$ pairs of endpoints, in an $\Omega(\ell)$-connected, or even $\ell^{\Omega(1)}$-connected graph. Our contributions in that direction can be of independent interest to the reader.
Theorem 2: To sketch our second main result, Theorem 2, we recall the proof of Gallai-Milgram theorem for undirected graphs. It is based on the following observation: if there are more than $\alpha(G)$ paths in a path cover of $G$, then there is an edge between two endpoints of distinct paths. Then these two paths can be joined into a single one and the size of the path cover is reduced by one. This proof gives a polynomial-time algorithm constructing a path cover of $G$ consisting of at most $\alpha(G)$ paths. Note that the algorithm does not have to know or compute $\alpha(G)$ itself: it is enough to iteratively join a pair of paths with adjacent endpoints until it is no longer possible. The algorithm can start from a trivial path cover consisting of $|V(G)|$ empty paths.
The procedure above on its own does not provide any additional relation between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ except for $p\leq\alpha(G)$. We propose a way to deal with this issue. First, we distinguish the paths in $\mathcal{P}$ into usual and special. We call a path usual if its endpoints are distinct and there is no edge between them in $G$. Otherwise, we call a path special. We argue that having usual paths in $\mathcal{P}$ is beneficial. For example, if all paths in $\mathcal{P}$ are usual, then the endpoints of paths in $\mathcal{P}$ form an independent set of size $2p$ as, otherwise, we would join two paths. In particular, this implies that $p\leq\alpha(G)/2$. Even having $k$ usual paths in $\mathcal{P}$ is enough since this yields an independent set of size $p+k$. Together with $\mathcal{P}$, this independent set is a feasible output for the algorithm of Theorem 2.
While special paths are unsuitable for constructing a large enough independent set, they are suitable for joining. A special path is either a path of length zero or one, or a cycle of length at least three (if considered together with the edge between the endpoints). Then, an edge between two arbitrary vertices of two special paths is as helpful as an edge between endpoints: it allows us to join the two paths and cover the vertices of their union by a single path. Our algorithm exploits this property of special paths and joins them exhaustively until there is no edge between special paths.
This property of special paths, in fact, gives us even more insight into the structure of $G$. Since special paths are independent of each other, and there are less than $k$ usual paths in $\mathcal{P}$, taking just one vertex from each special path constitutes an independent set of size greater than $p-k$. Taking two non-adjacent vertices instead of one from $2k$ special paths gives an independent set of size at least $p+k$, solving the problem immediately. Hence, the remaining case is when all but at most $2k$ special paths have vertex sets inducing cliques in $G$.
However, this structural insight alone is only half of the solution. To proceed further, we need another and more sophisticated way of transforming the paths in $\mathcal{P}$. Instead of joining two paths as before, this transformation takes two special paths and one usual path, so there are two edges without common endpoints connecting the special paths to the usual path. The transformation produces three paths, and at least two of these paths are usual. Hence, this transformation reduces the number of special paths and does not increase $\mathcal{P}$.
The transformation above provides the remaining structural insight for the algorithm. When this transformation and all the transformations above cannot be applied, a vertex set $S$ of size less than $k$ separates all usual paths from at least $p-4k$ special paths with vertex sets inducing cliques. In an auxiliary Lemma 4, we prove that the small size of $S$ can be exploited and irrelevant cliques can be removed from $G$ with respect to the size of an optimal path cover and the independence number. After applying this polynomial-time routine, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
The proof outline has come to the final case when $p=\mathcal{O}(k^{2})$. In this case, with Theorem 1 in hand, we aim to find the optimal path cover of $G$ as discussed previously in this section. The parameter choice in Theorem 1 is essential, and we choose it to be $p+k$. Thus, if the application of Theorem 1 does not give an optimal path cover, it gives an independent set of size $p+k$. Either way, the algorithm produces a solution feasible for Theorem 2.
We finish the outline of the proof of Theorem 2 by noting that the algorithm described above takes polynomial time, except for the application of Theorem 1.

### v2

@@ 3.2 The case of $\operatorname{pc}(G)<k^{C}$
Proof.
We present an algorithm $\mathcal{A}$. The algorithm assumes that $G$ contains at least one edge as, otherwise, $pc(G)=n$ and the minimum-size path cover consists of trivial paths.
Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$. To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$. On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching. Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial. The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 2 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$. The subroutine works in
running time.
If the outcome of the subroutine is an independent set $I$ in $G$ of size $p^{\prime}+k$, then the algorithm outputs $\mathcal{P}^{\prime}$ and $I$ as the solution. Since $|\mathcal{P}^{\prime}|=p^{\prime}$, $|I|\geq|\mathcal{P}^{\prime}|+k$ is satisfied, so this solution satisfies the lemma statement.
If the subroutine decides that $H_{ij}$ does not have TM-model in $G$ or the maximum TM-model of $H_{ij}$ in $G$ is not spanning, then $G$ cannot be covered with $i$ paths and $\mathcal{A}$ continues to the next iteration. If the subroutine gives a spanning TM-model of $H_{ij}$ in $G$, then the minimum path cover size of $G$ equals $i$ and $\mathcal{A}$ transforms this TM-model into a path cover of $G$ with $i$ paths. Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops.
Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$. The proof of the lemma is complete. ∎
3.3 Proof of Theorem 1
We are ready to prove Theorem 1. We describe an algorithm that starts from a trivial path cover of $G$ and then gradually transforms it. These transformations are indicated as reduction rules and are performed in polynomial time. We show that each such rule does not increase the number of paths in the path cover. Moreover, each reduction rule necessarily reduces either the total number of paths or the number of special (to be defined later) paths in the path cover.
A series of consecutive applications of the reduction rules always lead the algorithm to cases when the algorithm can produce a solution satisfying theorem statement in $f(k)\cdot n^{\mathcal{O}(1)}$ time and stop. We indicate such cases and the corresponding behavior of the algorithm as solution subroutines. The algorithm runtime thus always ends with a solution subroutine run.
In what follows, we present the reduction rules and solution subroutines. Each of the reduction rules reduces the number of paths in a path cover or reduces the number of some special paths. Solution subroutines output path covers and independent sets. The rules and subroutines are given in the order in which the algorithm checks their applicability. That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied.
When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties. In this case, the final solution subroutine calls the algorithms of Lemmas 1 and 2.
Reduction Rule 1.
The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one. Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$. When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
Special and usual paths. We call a path $P_{i}$ in the path cover special, if either $s_{i}=t_{i}$, or $P_{i}$ consists of only edge $s_{i}t_{i}$, or its endpoints are connected by the edge $s_{i}t_{i}$. If a path $P_{i}$ is not special, we call it usual. We claim that sufficiently many usual paths give a solution to the problem.
Claim 2.
Proof of Claim 2. We show that $G$ has an independent set $I$ of size $m+k$. Initialize $I$ with $\{s_{1},s_{2},\ldots,s_{m}\}$. Since 1 is not applicable, $I$ is independent in $G$.
Let $i_{1},i_{2},\ldots,i_{k}$ be $k$ distinct indices of usual paths. Since each $P_{i_{j}}$ is usual, $t_{i_{j}}\neq s_{i_{j}}$ and $s_{i_{j}}$ is not connected with $t_{i_{j}}$. Also, 1 is not applicable, so there is no edge between $t_{i_{j}}$ and any vertex in $I$, while $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}\}$ is also an independent set in $G$.
Put $t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}$ in $I$. We obtain an independent set of size $m+k$ in $G$, consequently $\alpha(G)-k\geq m$. ∎
2 yields the corresponding solution subroutine for the algorithm.
Solution Subroutine 1.
If there are at least $k$ usual paths among $\mathcal{P}=P_{1},P_{2},\ldots,P_{m}$, output $\mathcal{P}$ and $I$ as in 2.
Reduction Rule 2.
  1. 1.
    Add edge $s_{i}t_{i}$ to $P_{i}$, so it becomes a cycle;
  2. 2.
  3. 3.
    Join $P_{i}$ and $P_{j}$ via $us_{j}$ (or $ut_{j}$).
Reduction Rule 3.
  1. 1.
    Add edge $s_{j}t_{j}$ to $P_{j}$, so it becomes a cycle;
  2. 2.
    Remove edge $u_{j}v_{j}$ from $P_{j}$, so it becomes an $u_{j}v_{j}$-path.
  3. 3.
    Apply 2 to $P_{i}$, $P_{j}$ and $u_{i}u_{j}$.
Claim 3.
Proof of Claim 3. Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$. If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable. Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$. If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction. Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$. In this case, 1 is applicable, leading us to the final contradiction. ∎
The claim is the basis for the next solution subroutine, which the algorithm applies next. We explain how it works by proving the following claim.
Claim 4.
Proof of Claim 4. We construct an independent set $I$ consisting of more than $m+k$ vertices. Since 1 is not applicable, there are more than $m-k$ special paths. For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$. For each other special path, put its arbitrary vertex into $I$.
By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎
The solution subroutine based on 4 is given below. The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
Solution Subroutine 2.
If there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then output $\mathcal{P}$ and an $I$ as in 4.
Reduction Rule 4.
  1. 1.
  2. 2.
  3. 3.
    Join $P_{i}$ with the prefix part via $u_{i}x_{i}$;
  4. 4.
    Join $P_{j}$ with the suffix part via $u_{j}x_{j}$.
Claim 5.
Proof of Claim 5. The second part of the claim is trivial, since 4 produces three (or two, if $x_{i}x_{j}\in E(P_{\ell})$) paths by transforming three distinct paths $P_{i},P_{j},P_{\ell}$, so $|\mathcal{P}|$ cannot increase.
To prove the first part, we show that the resulting $v_{i}s_{\ell}$-path and $v_{j}t_{\ell}$-path are usual. We denote these two paths by $Q_{i}$ and $Q_{j}$ correspondingly. We give the proof only for $Q_{i}$, since the proof for $Q_{j}$ is symmetric.
We first show that $|V(Q_{i})|\geq 3$. Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$. The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly. Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold. But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$. Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$. This contradiction proves $|V(Q_{i})|\geq 3$.
It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$. Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable. This contradiction proves the claim. ∎
Claim 6.
If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$. Moreover, such a set $S$ could be constructed in polynomial time.
Proof of Claim 6. Since 2 is not applicable, there are less than $k$ usual paths in $\mathcal{P}$. We construct $S$ by taking one or zero vertices from each usual path, so $|S|<k$ holds automatically. We now explain the choice of the vertex to put in $S$ for each usual path.
Consider a usual path $P_{i}\in\mathcal{P}$. We call a vertex $u\in V(P_{i})$ a connector if there is a special path $P_{j}$ and a vertex $v\in V(P_{j})$ with $uv\in E(G)$. If $P_{i}$ has exactly one connector, then we put this connector in $S$. Otherwise, i.e. when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$.
We show that at least $m-2k$ special paths are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded.
Assume now that $P_{j}$ is a special path that is not isolated in $G-S$. Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$. By 3, $P_{i}$ can only be usual. Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors. As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$. Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$. Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$. ∎
Solution Subroutine 3.
Perform the following steps:
  1. 1.
    Obtain $S$ according to 6 and put $h:=\max\{0,m-4k\}$;
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
    Output $\mathcal{S}$ and (if obtained) $I$.
Claim 7.
If Reduction Rules 1-4 and Solution Subroutines 1,2 are not applicable, then 3 works in $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$ time and produces a feasible solution.
Proof of Claim 7. First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 6 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
We now move on to establish the upper bound on the running time. All steps, except for applying Lemma 2, are carried out in polynomial time. We have
By applying Lemma 2 to $G^{\prime},k^{\prime}$ and $\mathcal{P}^{\prime}$, we obtain the running time $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$.
Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$. Hence it is also a path cover. On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$.
Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$. First, let us consider the scenario where $I$ is not produced. This can occur only when the algorithm of Lemma 2 does not produce $I^{\prime}$. Then Lemma 2 guarantees that $\mathcal{S^{\prime}}$ is an optimal path cover of $G^{\prime}$ and $|\mathcal{S}^{\prime}|=\operatorname{pc}(G^{\prime})$. By Lemma 1, removal of $V(P_{j})$ for any $j\in[h]\setminus X$ decreases $\operatorname{pc}(G)$ by one, so $\operatorname{pc}(G^{\prime})=\operatorname{pc}(G)-(h-|X|).$ Since $|\mathcal{P}\setminus\mathcal{P}^{\prime}|=h-|X|$, we have $|\mathcal{S}|=\operatorname{pc}(G)$. It means that $\mathcal{S}$ is an optimal path cover of $G$.
Finally, let us consider the situation where $I$ is produced. In this case, the algorithm in Lemma 2 provides $\mathcal{S^{\prime}}$ and $I^{\prime}$ with $|I^{\prime}|-|\mathcal{S}^{\prime}|=k^{\prime}=2k$. We notice that
The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
This completes the proof of Theorem 1.